You bought a $200K condo. You got a 15-year fixed-rate mortgage and made a 20% down payment. What is your monthly payment if the annual interest rate is 4%?
Answer #1

value of the loan=200000*(1-20%)=160000

monthly payment

=PV/((1-(1+(r/m))^(n*m))/(n/m))

=160000/((1-(1+(4%/12))^(-15*12))/(4%/12))

=1183.50

the above is answer..